# HiMo: High-Speed Objects Motion Compensation in Point Clouds

### v2

@@ Abstract
Index Terms:
Range Sensing; Autonomous Driving Navigation; Computer Vision for Transportation; Motion Compensation
00footnotetext: 1Authors are with the Division of Robotics, Perception, and Learning (RPL), KTH Royal Institute of Technology, Stockholm 114 28, Sweden. (email: qingwen@kth.se)
2Authors are with Autonomous Transport Solutions Lab, Scania Group, Södertälje 151 87, Sweden.
Work done while interning at Scania.
@@ I Introduction
Light Detection and Ranging (LiDAR) sensors are an integral part of perception systems for autonomous driving. These sensors provide detailed depth information and point cloud data, complementing other sensing modalities (such as cameras) to offer high-precision 3D scene understanding [1, 2, 3] [4, 5, 6, 7]. However, due to the rotating mechanism of mechanical LiDAR sensors, different parts of the environment are measured at different times. This introduces motion-induced point cloud distortion in dynamic scenes, making it difficult to obtain accurate environment representations. We refer to this as rolling shutter distortion [8] to highlight the close relation with the effect seen in the image domain [9, 10].
@@ II-A Motion Compensation
The second component of this distortion, non-ego motion distortion, occurs due to the motion of other agents in the scene. To the best of our knowledge, no existing work addresses the correction of such distortion in raw point cloud data. While dynamic object segmentation methods [20, 21, 22] are related to motion understanding, they do not focus on distortion correction. The recent preprint SMORE [27] mentions this dynamic rolling shutter distortion in point clouds. However, they focus solely on how this distortion causes bad object mesh reconstruction and propose to improve the reconstructed meshes by combining multiple frames and output meshes. Instead, our work focuses on the distortion correction of raw data inside a single LiDAR frame, with many possible downstream applications. Additionally, the proposed HiMo pipeline is self-supervised, while SMORE requires ground truth tracking labels or a tracking network trained with annotated data.
@@ II-B Scene Flow Estimation
Scene flow estimation is the task of describing a 3D motion field between temporally successive point clouds [23, 28] [29, 30, 31, 32, 8]. Existing works applied to autonomous driving datasets can be categorized into supervised [24, 33] [34, 35] and self-supervised flow estimation [36, 37, 38, 39, 40, 41, 42, 43].
@@
To train models without labeled data or directly optimize at runtime, researchers propose self-supervised pipelines for scene flow estimation [36, 38, 39, 40, 47]. Neural Scene Flow Prior (NSFP) [38] provides high-quality scene flow estimates by optimizing ReLU and MLP layers at test time to minimize the Chamfer distance and maintain cycle consistency. FastNSF [39] leverages the same optimization but achieves significant speedup by computing the Chamfer loss using distance transform [48]. ICP-Flow [40] performs Iterative Closest Point (ICP) between each cluster in two point clouds and trains a feedforward neural network for real-time inference. SeFlow [36] integrates dynamic awareness and proposes four novel self-supervised loss terms to allow for efficient training using large datasets.
Existing scene flow estimation methods are not intended to address non-ego motion distortions in raw point cloud data. However, they capture point-level velocity information that is valuable for the task. Guided by this insight, in this work, we repurpose scene flow estimation as a key component to correct non-ego motion distortions. In developing HiMo, we observed that existing self-supervised scene flow estimators often produce degenerate or inconsistent estimates under high-speed distortion. To mitigate these inconsistencies, we introduce SeFlow++, a scene flow estimator built on SeFlow but with higher training efficiency and better performance, particularly in high-speed scenarios. Specifically, we design a dynamic auto-labelling module (See Section IV-A) to improve the self-supervised signal at high-speed motion and a symmetric loss computation (See Section IV-B) to improve training efficiency with limited data.
Figure 3: Top-view example of LiDAR sweeps showing how distortions are created for vehicles with (a) a single LiDAR and (b) two LiDARs. The small concentric circles of yellow and green are LiDARs. Both cases cause a displacement distance for the high-speed object. Light and dark cars show one car’s movement between two timestamps. (a) One complete single LiDAR scan sweep (small yellow dots are the first and last scan points). This case is only observed when moving objects are at the scan boundaries. (b) Two LiDAR scans separated in orientation by 180 degrees. This case is always observed for fast-moving objects. An animation illustrating both cases is included in the supplementary video.
@@ III-B HiMo Pipeline
$\mathcal{D}(\mathbf{p})=\mathcal{V}(\mathbf{p})\Delta T(\mathbf{p}).$ (1)
@@ IV Scene Flow
Scene flow is the core module in our HiMo motion compensation pipeline. To train or optimize a scene flow estimator, we need a supervision signal and a proper loss function. In supervised training, the signal is provided by human-labeled ground truth flow [33, 35]. However, such human annotation is costly and potentially error-prone due to the distortions of high-speed vehicles in raw LiDAR data. We therefore focus on self-supervised methods that do not rely on annotations. In our experiments, we found that existing self-supervised scene flow methods either require huge computational resources or do not perform well under the conditions of scarce training data and high-speed object distortions. To adapt to the high-speed regime and reduce the requirement on the amount of training data, we present SeFlow++, a new scene flow method based on the previous state-of-the-art method SeFlow [36]. Compared to SeFlow, SeFlow++ refines the dynamic auto-labeler to provide more robust supervision signals and incorporates symmetric loss functions for improved training efficiency and better temporal consistency.
@@ IV-A Auto-labeler
To obtain a self-supervision signal, SeFlow [36] classifies points into static and dynamic based on DUFOMap output [49]. The dynamic points are then grouped into clusters, as shown in Fig. 6 (top). However, we observe that the scene flow estimator trained using the SeFlow strategy is not always accurate. This is because errors in the dynamic point classification of DUFOMap get propagated into the subsequent object clustering step, resulting in misclassification of dynamic and static points. To refine the auto-labeling in SeFlow++, we compute the dynamic clusters by aggregating more information. As shown in Fig. 6 (bottom), this process starts with the parallel computation of clusters and point-level dynamic classification:
@@ IV-A2 Dynamic classification
$\mathcal{P}_{\text{nnd}}=\{\mathbf{p}|\mathrm{D}_{\text{min}}(\mathbf{p},\mathcal{P}_{t+1})>\mathrm{D}_{thres},\mathbf{p}\in\mathcal{P}_{t}\},$ (3)
where $\mathrm{D}_{\text{min}}(\mathbf{p},\mathcal{P}_{t+1})$ represents the distance between point $\mathbf{p}$ and its nearest neighbor in $\mathcal{P}_{t+1}$, and $D_{thres}$ is a user defined threshold.
@@
where $\mathcal{C}$ is the set of all clusters based on HDBSCAN on $\mathcal{P}$, $c$ and $\mathbf{p}$ represent an individual cluster and a point in the cluster, respectively. The function $f$ integrates the two dynamic labels $\mathcal{P}_{\text{dufo}}$ and $\mathcal{P}_{\text{nnd}}$ and is defined as follows:
$f(c)=\begin{cases}\tt{true},&\text{if }\min(r_{1},r_{2})\geq\tau_{1}~{}\&~{}\max(r_{1},r_{2})\geq\tau_{2}\\ \tt{false},&\text{otherwise}\end{cases}$ (5)
where $r_{1}=\frac{|\mathcal{P}_{\text{dufo}}(c)|}{|\mathcal{P}_{c}|}$ and $r_{2}=\frac{|\mathcal{P}_{\text{nnd}}(c)|}{|\mathcal{P}_{c}|}$ represent the proportion of dynamic points labeled in the cluster instance $c$ by the two methods respectively. $|\mathcal{P}|$ denotes the cardinality of (i.e., the number of points in) the point cloud $\mathcal{P}$. Conceptually, this means that we examine each individual object clusters $c$ separately, and assign all points inside this cluster as dynamic only if the propotion of dynamic points in $c$ according to $\mathcal{P}_{\text{dufo}}$ exceeds $\tau_{1}$, and the proportion according to $\mathcal{P}_{\text{nnd}}(c)$ exceeds $\tau_{2}$. The decision thresholds $\tau_{1},\tau_{2}$ are hyperparameters. In all our experiments, these were set to 5% and 10%, respectively.
@@ IV-B Self-supervised Loss
Although SeFlow achieves state-of-the-art performance in self-supervised scene flow estimation, we noticed that it occasionally produces degenerate or inconsistent flow estimation between consecutive frames. This inconsistency introduces errors when we apply the estimated flow to correct motion distortions. Prior works have introduced cycle consistency to improve scene flow estimation and mitigate degenerate solutions, particularly in sparse point clouds [52, 38]. However, these methods rely on explicit backward flow computation or duplicated networks, which are inefficient. To encourage consistent flow estimate with minimal overhead, we propose a symmetric flow loss that can be calculated within a single forward pass. Note that due to the symmetry around $\mathcal{P}_{t}$, applying the forward flow with a negative sign (--) effectively simulates a backward flow from $\mathcal{P}_{t}$ to $\mathcal{P}_{t-1}$. Guided by this insight, we reformulate the first chamfer distance loss in Eq. 7 to a symmetric version as follows:
@@
$\mathcal{L}_{\text{static}}=\frac{1}{|{\mathcal{P}}_{s}|}\sum_{\mathbf{p}\in\mathcal{P}_{s}}||\Delta\hat{\mathcal{F}}(\mathbf{p})||_{2}^{2},$ (9)
@@
$\mathcal{L}_{\text{dcls}}=\frac{1}{|\mathcal{P}_{t,d}|}{\sum_{c_{i}\in\mathcal{C}_{t,d}}\left(\sum_{p_{j}\in\mathcal{P}_{c_{i}}}||\hat{f}_{p_{j}}-\tilde{f}_{c_{i}}||^{2}_{2}\right)}.$ (12)
@@ V-A Dataset
More interactive visualization results on Argoverse 2 [17] and ZOD [26] are available on the project page111 https://kin-zhang.github.io/HiMo .
@@ V-B Annotated Ground truth
TABLE I: Object motion compensation result comparisons using different scene flow methods in our HiMo pipeline on the Scania validation set. The first row reports the errors for raw data with ego-motion compensation only, and the rest is our HiMo pipeline motion compensation with different scene flow estimators as an ablation study. Upper groups are supervised methods with scene flow networks trained on Argoverse 2 [17] and inference directly on the Scania dataset. Lower groups are self-supervised methods. $\pm$ means the standard deviation in the evaluation data (100 frames). All methods in the HiMo pipeline achieve better accuracy performance and object shape similarity compared to raw data. Our proposed SeFlow++ achieves state-of-art compensated performance in shape description for both car and other vehicle object types. The ↓blue value represents the error percentage decrease relative to the distorted raw data.
@@ V-C Evaluation Metrics
where ||$\cdot$|| ⋅ | denotes cardinality of a set, $\mathcal{C}_{gt}$ denotes the ground truth clusters in the frame, $\mathcal{P}_{c_{i}}$ and $\mathcal{P}_{\mathcal{C}_{gt}}$ denote the set of points in cluster $c_{i}$ and in all ground truth clusters, $\mathcal{P^{\prime}}_{est,c_{i}}$ and $\mathcal{P^{\prime}}_{gt,c_{i}}$ denote the point set $c_{i}$ compensated using the estimated motion and ground truth, respectively. The smaller the CDE, the greater the shape similarity between the estimated instance shape and the ground truth.
While we believe that shape similarity best measures a method’s ability to undistort moving objects, its reliance on Chamfer distance implies that it is based on nearest neighbor matching. Such maching results do not guarantee correct association at point level. Therefore, we also compute the point-level accuracy that can be represented as the mean point error (MPE),
@@ V-D Evaluated Methods
To address non-ego motion distortion compensation, we propose the HiMo pipeline (see Fig. 5), which leverages scene flow estimation to perform motion compensation. Given consecutive frames of point cloud data, HiMo employs any scene flow estimators to estimate point velocities and then transform these velocities into distortion distances. These distortion distances then can be used to undistort all motions in the raw point cloud.
As mentioned in Section III-B, the HiMo pipeline is agnostic to the choice of scene flow estimators. To evaluate the performance and highlight the flexibility of HiMo, we incorporate different state-of-the-art scene flow estimators into the pipeline. The evaluated scene flow estimators [^2] The scene flow performance of these scene flow methods are provided in Section -B. are outlined below:
@@
    SeFlow++ (Ours): Proposed in this paper, it improves SeFlow with refined dynamic auto-labeling (see Section IV-A) and symmetric forward flow self-supervise loss supervision (see Section IV-B).
@@
All code to reproduce results and run the HiMo pipeline can be found in https://github.com/KTH-RPL/HiMo. The main training hyperparameters are listed here: learning rate ($2e^{-4}$) with Adam optimizer [55], batch size ($8$), the total training epoch ($9$). More configurations can be found in the code. All experiments are executed on a desktop powered by an Intel Core i9-12900KF CPU and equipped with a GeForce RTX 3090 GPU.
@@ VI-A Quantitative Results
The comparative analysis of different scene flow methods using the HiMo pipeline on the Scania dataset is detailed in Table I. The baseline results from Ego-motion Compensation are shown on the first row, with large values on both CDE and MPE. Perfect undistortion would result in both metrics being zero. As shown in the rest of Table I, compared to this baseline, the HiMo pipeline reduces distortion errors in both CDE and MPE regardless of the scene flow estimator it is coupled with. This demonstrates the effectiveness of our HiMo pipeline in motion compensation, with up to 81% in shape improvement. However, when combined with different scene flow estimators, the performance of the HiMo pipeline differs. Despite not having seen Scania data before, DeFlow performs well on the CAR category according to both CDE and MPE. This is because the car-type objects in Argoverse 2 and Scania data are similar in shape. Comparatively, DeFlow does not perform as well on the OTHERS category. This is because the OTHERS category contains objects with high motion speed and long vehicle sizes, with heavier motion-related distortions in raw data. Such distortions cause difficulty in transferring the knowledge from the previous dataset. Compared to supervised methods, most self-supervised methods achieve both lower CDE and MPE on the OTHERS category. The only exception is ICP-Flow, which performs on par or worse than the supervised methods on both CAR and OTHERS category. This is because ICP-Flow employs many heuristics in its optimization and ICP matching procedure. Hence, extensive parameter tweaking is needed for different data and scenarios. As shown in Table II, ICP-Flow performs much better on Argoverse 2, the dataset it is optimized for. The best performance in shape similarity, quantified by the lowest CDE, is achieved by the HiMo pipeline with our proposed SeFlow++. This setup also shows competitive performance on MPE.

### v3

@@ I Introduction
Light Detection and Ranging (LiDAR) sensors are an integral part of perception systems for autonomous driving. These sensors provide detailed depth information and point cloud data, complementing other sensing modalities (such as cameras) to offer high-precision 3D scene understanding [1, 2, 3, 4, 5, 6, 7]. However, due to the rotating mechanism of mechanical LiDAR sensors, different parts of the environment are measured at different times. This introduces motion-induced point cloud distortion in dynamic scenes, making it difficult to obtain accurate environment representations. We refer to this as rolling shutter distortion [8] to highlight the close relation with the effect seen in the image domain [9, 10].
@@ II-A Motion Compensation
The second component of this distortion, non-ego motion distortion, occurs due to the motion of other agents in the scene. To the best of our knowledge, no existing work addresses the correction of such distortion in raw point cloud data. While dynamic object segmentation methods [20, 21, 22] are related to motion understanding, they do not focus on distortion correction. The recent work SMORE [27] mentions this dynamic rolling shutter distortion in point clouds. However, their work focuses on combining multiple frames of distorted data to improve the quality of reconstructed object meshes, without correcting the raw data itself. Instead, our work focuses on the distortion correction of raw data inside a single LiDAR frame, with many possible downstream applications. Additionally, the proposed HiMo pipeline is self-supervised, while SMORE requires ground truth tracking labels or a tracking network trained with annotated data.
@@ II-B Scene Flow Estimation
Scene flow estimation is the task of describing a 3D motion field between temporally successive point clouds [23, 28, 29, 30, 31, 32, 8]. Existing works applied to autonomous driving datasets can be categorized into supervised [24, 33, 34, 35] and self-supervised flow estimation [36, 37, 38, 39, 40, 41, 42, 43].
@@
To train models without labeled data or directly optimize at runtime, researchers propose self-supervised pipelines for scene flow estimation [36, 38, 39, 40, 47]. Neural Scene Flow Prior (NSFP) [38] provides high-quality scene flow estimates by optimizing MLP layers at test time to minimize the Chamfer distance and maintain cycle consistency. FastNSF [39] leverages the same optimization but achieves significant speedup by computing the Chamfer loss using distance transform [48]. ICP-Flow [40] performs Iterative Closest Point (ICP) between each cluster in two point clouds and trains a feedforward neural network for real-time inference. SeFlow [36] integrates dynamic awareness and proposes novel self-supervised loss terms to allow for efficient training using large datasets.
Existing scene flow estimation methods are not intended to address non-ego motion distortions in raw point cloud data. However, they capture point-level velocity information that is valuable for the task. Guided by this insight, in this work, we repurpose scene flow estimation as a key component to correct non-ego motion distortions. In developing HiMo, we observed that existing self-supervised scene flow estimators often produce degenerate or inconsistent estimates under high-speed distortion. To mitigate these inconsistencies, we introduce SeFlow++, a scene flow estimator built on SeFlow but with higher training efficiency and better performance, particularly in high-speed scenarios. Specifically, we design a dynamic auto-labelling module (See Section IV-A) to improve the self-supervised signal at high-speed motion and a symmetric loss computation (See Section IV-B) with a larger three-frame model backbone to improve training efficiency with limited data.
Figure 3: Top-view example of LiDAR sweeps showing how distortions are created for vehicles with (a) a single LiDAR and (b) two LiDARs. The small concentric circles of yellow and green are LiDARs. Both cases cause a displacement distance for the high-speed object. The light and dark cars show the vehicle’s positions at two timestamps (the beginning and the end of the LidDAR sweep). (a) One complete single LiDAR scan sweep (small yellow dots are the first and last scan points). This case is only observed when moving objects are at the scan boundaries. (b) Two LiDAR scans separated in orientation by 180 degrees. This case is always observed for fast-moving objects. An animation illustrating both cases is included in the supplementary video.
@@ III-B HiMo Pipeline
$\mathcal{D}(\mathbf{p})=\mathcal{V}(\mathbf{p})\Delta T(\mathbf{p}),$ (1)
@@ IV Scene Flow
Scene flow is the core module in our HiMo motion compensation pipeline. To train or optimize a scene flow estimator, we need a supervision signal and a proper loss function. In supervised training, the signal is provided by human-labeled ground truth flow [33, 35]. However, such human annotation is costly and potentially error-prone due to the distortions of high-speed vehicles in raw LiDAR data. We therefore focus on self-supervised methods that do not rely on annotations. In our experiments, we found that existing self-supervised scene flow methods either require huge computational resources or do not perform well under the conditions of scarce training data and high-speed object distortions. To adapt to the high-speed regime and reduce the requirement on the amount of training data, we present SeFlow++, a new scene flow method based on the previous state-of-the-art method SeFlow [36]. Compared to SeFlow, SeFlow++ enhances training efficiency and temporal consistency by combining symmetric loss functions with a larger three-frame model backbone, and additionally provides more robust supervision signals through a refined dynamic auto-labeler.
@@ IV-A Auto-labeler
To obtain a self-supervision signal, SeFlow [36] classifies points into static and dynamic based on DUFOMap [49]. The dynamic points are then grouped into clusters, as shown in Fig. 6 (top). However, we observe that the scene flow estimator trained using the SeFlow strategy is not always accurate. This is because errors in the dynamic point classification of DUFOMap get propagated into the subsequent object clustering step, resulting in misclassification of dynamic and static points. To refine the auto-labeling in SeFlow++, we compute the dynamic clusters by aggregating more information. As shown in Fig. 6 (bottom), this process starts with the parallel computation of clusters and point-level dynamic classification:
@@ IV-A2 Dynamic classification
$\mathcal{P}_{\text{nnd}}=\{\mathbf{p}|\mathrm{D}_{\text{min}}(\mathbf{p},\mathcal{P}_{t+1})>\tau_{d},\mathbf{p}\in\mathcal{P}_{t}\},$ (3)
where $\mathrm{D}_{\text{min}}(\mathbf{p},\mathcal{P}_{t+1})$ represents the distance between point $\mathbf{p}$ and its nearest neighbor in $\mathcal{P}_{t+1}$, and $\tau_{d}$ is a user defined threshold.
@@
where $\mathcal{C}$ is the set of all clusters based on HDBSCAN on $\mathcal{P}$, $c$ and $\mathbf{p}$ represent an individual cluster and a point in the cluster, respectively. The function $f$ integrates the two dynamic labels $\mathcal{P}_{\text{dufo}}$ and $\mathcal{P}_{\text{nnd}}$ is defined as follows:
$f(c)=\begin{cases}\tt{true},&\text{if }\min(r_{1},r_{2})\geq\tau_{1}~\&~\max(r_{1},r_{2})\geq\tau_{2}\\ \tt{false},&\text{otherwise}\end{cases},$ (5)
where $r_{1}=\frac{|\mathcal{P}_{\text{dufo}}(c)|}{|\mathcal{P}_{c}|}$ and $r_{2}=\frac{|\mathcal{P}_{\text{nnd}}(c)|}{|\mathcal{P}_{c}|}$ represent the proportion of dynamic points labeled in the cluster instance $c$ by the two methods respectively. $|\mathcal{P}|$ denotes the cardinality of (i.e., the number of points in) the point cloud $\mathcal{P}$. Conceptually, this means that we examine each individual object clusters $c$ separately, and assign all points inside this cluster as dynamic only if both methods identify a proportion of dynamic points of at least $\tau_{1}$, and at least one of them provides a stronger signal with a proportion of at least $\tau_{2}$. The decision thresholds $\tau_{1},\tau_{2}$ are hyperparameters.
@@ IV-B Self-supervised Loss
Although SeFlow achieves state-of-the-art performance in self-supervised scene flow estimation, we noticed that it occasionally produces degenerate or inconsistent flow estimation between consecutive frames. This inconsistency introduces errors when we apply the estimated flow to correct motion distortions. Prior works have introduced cycle consistency to improve scene flow estimation and mitigate degenerate solutions, particularly in sparse point clouds [52, 38]. However, these methods rely on explicit backward flow computation or duplicated networks, which are inefficient. To encourage consistent flow estimate with minimal overhead, we propose a symmetric flow loss that can be calculated within a single forward pass. Note that due to the symmetry around $\mathcal{P}_{t}$, applying the forward flow with a negative sign (-) effectively simulates a backward flow from $\mathcal{P}_{t}$ to $\mathcal{P}_{t-1}$. Guided by this insight, we reformulate the first chamfer distance loss in Eq. 7 to a symmetric version as follows:
@@
[…]
@@
[…]
@@ V-A Dataset
More interactive visualization results on Argoverse 2 [17] and ZOD [26] are available on the project page [^1] https://kin-zhang.github.io/HiMo.
@@ V-B Annotated Ground truth
TABLE I: Object motion compensation result comparisons using different scene flow methods in our HiMo pipeline on the Scania validation set. The first row reports the errors for raw data with ego-motion compensation only, and the rest is our HiMo pipeline motion compensation with different scene flow estimators as an ablation study. Upper groups are supervised methods with scene flow networks trained on Argoverse 2 [17] and inference directly on the Scania dataset. Lower groups are self-supervised methods. $\pm$ means the standard deviation in the evaluation data (100 frames). All methods in the HiMo pipeline achieve better accuracy performance and object shape similarity compared to raw data. Our proposed SeFlow++ achieves state-of-the-art compensated performance in shape description for both car and other vehicle object types. The ↓blue value represents the error percentage decrease relative to the distorted raw data.
@@ V-C Evaluation Metrics
where $|\cdot|$ denotes cardinality of a set, $\mathcal{C}_{gt}$ denotes the ground truth clusters in the frame, $\mathcal{P}_{c_{i}}$ and $\mathcal{P}_{\mathcal{C}_{gt}}$ denote the set of points in cluster $c_{i}$ and in all ground truth clusters, $\mathcal{P^{\prime}}_{est,c_{i}}$ and $\mathcal{P^{\prime}}_{gt,c_{i}}$ denote the point set $c_{i}$ compensated using the estimated motion and ground truth, respectively. The smaller the CDE, the greater the shape similarity between the estimated instance shape and the ground truth.
While we believe that shape similarity best measures a method’s ability to undistort moving objects, its reliance on Chamfer distance implies that it is based on nearest neighbor matching. Such matching results do not guarantee correct association at point level. Therefore, we also compute the point-level accuracy that can be represented as the mean point error (MPE),
@@ V-D Evaluated Methods
To address non-ego motion distortion compensation, we propose the HiMo pipeline (see Fig. 5), which leverages scene flow estimation to perform motion compensation. Given consecutive frames of point cloud data, HiMo employs any scene flow estimators to estimate point velocities and then transforms these velocities into distortion distances. These distortion distances then can be used to undistort all motions in the raw point cloud.
As mentioned in Section III-B, the HiMo pipeline is agnostic to the choice of scene flow estimators. To evaluate the performance and highlight the flexibility of HiMo, we incorporate different state-of-the-art scene flow estimators into the pipeline. The evaluated scene flow estimators [^2] The scene flow performance of these scene flow methods is provided in Section -B. are outlined below:
@@
    SeFlow++ (Ours): Proposed in this paper, it improves SeFlow with refined dynamic auto-labeling (see Section IV-A). It leverages a three-frame DeFlow backbone to implement a symmetric self-supervised loss (see Section IV-B), enhancing flow consistency and training efficiency.
@@
All code to reproduce results and run the HiMo pipeline can be found in https://github.com/KTH-RPL/HiMo. The main training hyperparameters are listed here for Scania highway datasets: learning rate ($2e^{-4}$) with Adam optimizer [55], batch size ($8$), the total training epoch ($9$). More configuration details can be found in the code. All runtime experiments are executed on a desktop powered by an Intel Core i9-12900KF CPU and equipped with a GeForce RTX 3090 GPU.
@@ VI-A Quantitative Results
The comparative analysis of different scene flow methods using the HiMo pipeline on the Scania dataset is detailed in Table I. The baseline results from Ego-motion Compensation are shown on the first row, with large values on both CDE and MPE. Perfect undistortion would result in both metrics being zero. As shown in the rest of Table I, compared to this baseline, the HiMo pipeline reduces distortion errors in both CDE and MPE regardless of the scene flow estimator it is coupled with. This demonstrates the effectiveness of our HiMo pipeline in motion compensation, with up to 81% in shape improvement. However, when combined with different scene flow estimators, the performance of the HiMo pipeline differs. Despite not having seen Scania data before, DeFlow performs well on the CAR category according to both CDE and MPE. This is because the car-type objects in Argoverse 2 and Scania data are similar in shape. Comparatively, DeFlow does not perform as well on the OTHERS category. This is because the OTHERS category contains objects with high motion speed and long vehicle sizes, with heavier motion-related distortions in raw data. Such distortions cause difficulty in transferring the knowledge from the previous dataset. Compared to supervised methods, most self-supervised methods achieve both lower CDE and MPE on the OTHERS category. The only exception is ICP-Flow, which performs on par or worse than the supervised methods on both CAR and OTHERS categories. This is because ICP-Flow employs many heuristics in its optimization and ICP matching procedure. Hence, extensive parameter tweaking is needed for different data and scenarios. As shown in Table II, ICP-Flow performs much better on Argoverse 2, the dataset it is optimized for. The best performance in shape similarity, quantified by the lowest CDE, is achieved by the HiMo pipeline with our proposed SeFlow++. This setup also shows competitive performance on MPE.